OpenStudy (liizzyliizz):

explain why the function is discontinuous at the given point ( f(x) = { x^2-2x-8/ x-4 if x isnt equal to 4 and 3 if x=4 .. ( is it because there is a hole at 4 ?)

OpenStudy (zarkon):

\[f(4)=3\] but \[\lim_{x\to 4}\frac{x^2-2x-8}{x-4}=6\]

OpenStudy (aravindg):

satellite help me amistre could not answer my question

OpenStudy (anonymous):

it isn't equal to anything at 4 because you will have \[4-4=0\] in the denominator

OpenStudy (zarkon):

she has a piecewise function

OpenStudy (aravindg):

satellite help me amistre could not answer my question

OpenStudy (aravindg):

satellite help me amistre could not answer my question

OpenStudy (anonymous):

it is true you can rewrite by factoring and canceling and gettin \[f(x)=x+2\] but that is only if \[x\neq4\]

OpenStudy (anonymous):

oh i should learn to read hold on

OpenStudy (liizzyliizz):

yeah it is x\[\neq4\] lol

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} \frac{x^2-2x-8}{ x-4} & \text{if} & x \neq 4 \\ 3& \text{if} & x = 4 \end{array} \right. \]

OpenStudy (liizzyliizz):

yess

OpenStudy (anonymous):

ooooh ok

OpenStudy (anonymous):

so \[\frac{x^2-2x-8}{x-4}=\frac{(x-4)(x+2)}{x-4}=x+2\] which holds if \[x\neq 4\]

OpenStudy (anonymous):

so this function, even though it looks like a rational function, is just the line \[y=x+2\] with a hole in it where (4,6) should be right?

OpenStudy (aravindg):

satellite help me after this???

OpenStudy (anonymous):

sure hold on

OpenStudy (liizzyliizz):

so i was right it was just the hole..

OpenStudy (anonymous):

meaning that if you wanted to "remove" the discontinuity, you would just call it \[f(x)=x+2\] and be done

OpenStudy (anonymous):

or if you want to spend a lot of ink you could write \[f(x) = \left\{\begin{array}{rcc} \frac{x^2-2x-8}{ x-4} & \text{if} & x \neq 4 \\ 6 & \text{if} & x = 4 \end{array} \right.\]

OpenStudy (anonymous):

because the function really wants to be 6 when x = 4. but some idiot decided to call it 3 there, so it is not continuous

OpenStudy (liizzyliizz):

lol the idiot would be my math book -_-

OpenStudy (anonymous):

and yes, you are right, it is a "hole" which is sometimes called a 'removable discontinuity" because if you define the function correctly you an "remove" it

OpenStudy (liizzyliizz):

yeah. ahh ok one last question my next problems involve using theorems (4,5,7, and 9) im not quite sure what that is.. for some functions.. x_x is it something my hw didnt specify?

OpenStudy (aravindg):

come